OpenStudy (anonymous):

what is one exterior angle of a regular 13-gon as a fraction??

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know what the interior angle is of a regular "n-gon" ?

OpenStudy (misty1212):

if not, say so and i will tell you wow, look at all the ghosts ! i see ten of them

OpenStudy (anonymous):

I know the formula is n-2(180)

OpenStudy (anonymous):

I need it as a fraction and I am very confused

OpenStudy (misty1212):

the formula for interior angles is \((n-2)\times \frac{180}{n}\) and for external angles (what you are looking for) is \(\frac{360}{n}\)

OpenStudy (anonymous):

ok so I do 360/13?

OpenStudy (misty1212):

your formula was for the SUM of the interior angles, not the measure of each angle yes, that is what you do \(\frac{360}{13}\)

OpenStudy (anonymous):

ok thank you so much!

OpenStudy (misty1212):

\(\color\magenta\heartsuit\)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

are you confused?

OpenStudy (misty1212):

me? i am always a bit confused. why do you ask?

OpenStudy (misty1212):

13 ghosts wasn't that a movie?